### v2

# eprint Nisho-1-2025

Abstract
We propose a new method for detecting dark matter axions using a resonant cavity coupled with a quantum Hall system. When a sample exhibiting quantum Hall effect is placed inside the cavity and the cavity is tuned to resonance, two-dimensional electrons absorb the amplified radiation, leading to a rise in the sample’s temperature. By monitoring this temperature increase, the mass $m_{a}$ of the axion can be inferred. As an example, consider a GaAs sample with small thickness $d=0.1\,\mu\mathrm{m}$ and its heat capacity $C_{s}$ at temperature $T=20\,\mathrm{mK}$. According to the QCD axion model, because the energy flux of the incoming radiation is $P_{ra}\sim 10^{-18}g_{\gamma}^{2}\mbox{W cm}^{-2}(10^{-5}\mbox{eV}/m_{a})(B/1.5\times 10^{5}\mbox{Gauss})^{2}(\rho_{d}/0.3\rm GeVcm^{-3})$ at the resonance with $g_{\gamma}(\rm KSVZ)=-0.96$ and $g_{\gamma}(\rm DFSZ)=0.37$, the temperature increase $(P_{ra}/C_{s})\times\tau$ is approximately $\simeq 0.25(\mbox{mK})(\tau/10\mbox{ms})\,g_{\gamma}^{2}\,(20\mbox{mK}/T)^{3}(10^{-5}\mbox{eV}/m_{a})(B/1.5\times 10^{5}\mbox{Gauss})^{2}(0.1\mu\mbox{m}/d)$, where $\tau$ is the time constant associated with the heat dissipation into thermal bath.
I introduction
Elucidating the nature of dark matter in the universe remains one of the most pressing challenges in modern physics. Among the various candidates, the axionaxion1 ; axion2 ; axion3 is of particular interest as it not only serves as a compelling dark matter candidate but also offers a potential solution to the strong CP problem in quantum chromodynamics (QCD). Despite numerous experimental efforts ( see the references in review ) to detect axions, they have yet to be observed.
In this paper, we propose a novel method for detecting axion dark matter using a resonant cavity coupled with the quantum Hall effect girvin . While resonant cavities have been employed in axion detection experiments under strong magnetic fields sikivie ; admx , detecting axions with masses $m_{a}$ greater than $10^{-5}$eV remains challenging due to the low sensitivity caused by the small volume of the cavity. Our approach overcomes this limitation.
By tuning the cavity to an appropriate scale, microwave signals generated by axions can be resonantly amplified. When a sample exhibiting the quantum Hall effect is placed inside the cavity, the two-dimensional (2D) electrons in the quantum Hall state absorb the amplified microwave radiation, resulting in an increase in temperature. This temperature rise serves as an observable signature of axion presence and allows for the determination of the axion mass.
( In previous papersiwaqhe1 ; iwaqhe2 ; iwaqhe3 we have pointed out that among the experiments conducted so far on the quantum Hall effect, there exist phenomena that appear to be influenced by dark matter axions. These phenomena are thought to result from axions converting into electromagnetic waves in a strong magnetic field, which are then absorbed into the quantum Hall state. )
The temperature increase due to energy absorption depends on the heat capacity of the semiconductor sample. The total heat capacity $C$ is the sum of the semiconductor heat capacity $C_{s}$ and the 2D electron contribution $C_{e}$. By appropriately tuning the Fermi energy, the contribution from the 2D electrons $C_{e}$ can be made negligible compared to $_{s}; CC$$\simeq C_{s}$. At low temperatures (e.g., 1010mK), the semiconductor heat capacity $C_{s}$ is dominated by phonons and follows the relation $C_{s}\propto T^{3}$, making the temperature increase highly sensitive to the temperature $T$. The lower the temperature, the greater the temperature rise for a given amount of absorbed energy. In our analysis, we assume a temperature of T=20T=20mK, which is relatively easy to achieve experimentally.
We show that the microwave power $P_{a}$ generated by axions within the cavity is absorbed by the 2D electron system. This results in a temperature rise at a rate given by $P_{ra}/C_{s}$, where $P_{ra}$ is the portion of the axion-induced power actually absorbed by the 2D electrons. Without resonance, this absorbed power is negligibly small, but at resonance, the temperature increase becomes detectable.
As a concrete example, for a GaAs sample involving the quantum Hall state, the temperature increase rate at resonance is approximately
$\frac{P_{ra}}{C}\simeq\frac{0.025\,\mathrm{mK}}{\mathrm{ms}}\times g_{\gamma}^{2}\left(\frac{20\,\mathrm{mK}}{T}\right)^{3}\Big(\frac{10^{-5}\mbox{eV}}{m_{a}}\Big)\left(\frac{B_{t}}{1.5\times 10^{5}\,\mathrm{Gauss}}\right)^{2}\left(\frac{0.1\,\mu\mathrm{m}}{d}\right)\Big(\frac{\rho_{d}}{0.3\rm GeVcm^{-3}}\Big),$ (1)
with the density $\rho_{d}$ of the dark matter axion, where $d$ is the sample thickness. $g_{\gamma}$ denotes QCD axion model; $g_{\gamma}(\rm KSVZ)\simeq-0.96$ and $g_{\gamma}(\rm DFSZ)\simeq 0.37$. The 2D electron layer is tilted with respect to the magnetic field $\vec{B}_{t}$, with an angle $\theta=\pi/6$ between the layer and the plane perpendicular to $\vec{B}_{t}$ (see Fig.2). This tilting makes possible absorption of the axion-induced microwave radiation.
It is important to note that this temperature rise rate is independent of the surface area $S$ of the sample, since both the absorbed power $P_{ra}\propto S$ and the heat capacity $C_{s}\propto S$. The form of semiconductor sample is rectangular with surface area $S$ and thickness $d$.
Finally, because the absorbed energy dissipates into the thermal bath, the actual observed temperature increase $\Delta T$ depends on the time constant $\tau$ of the dissipation. For example, with the time constant $\tau=10$ms, we have
$\Delta T\simeq 0.25\,\mathrm{mK}\times\left(\frac{\tau}{10\,\mathrm{ms}}\right)g_{\gamma}^{2}\left(\frac{20\,\mathrm{mK}}{T}\right)^{3}\Big(\frac{10^{-5}\mbox{eV}}{m_{a}}\Big)\Big(\frac{B_{t}}{1.5\times 10^{5}\mbox{Gauss}}\Big)^{2}\left(\frac{0.1\,\mu\mathrm{m}}{d}\right)\Big(\frac{\rho_{d}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
or for general coupling constant $g_{a\gamma\gamma}$,
$\Delta T\simeq 0.5\text{mK}\left(\frac{\tau}{10\,\text{ms}}\right)\,\Big(\frac{g_{a\gamma\gamma}}{10^{-14}\text{GeV}^{-1}}\Big)^{2}\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)^{3}\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$ (3)
It is favorable to cover the semiconductor sample with a transparent thin sheet possessing much small heat capacity and much small heat conductance for $\tau$ to be larger than 1010ms.
We begin by reviewing the fundamental properties of the quantum Hall state in 2D electron systems. We then demonstrate how microwave radiation, generated by axions in the presence of an external magnetic field $B_{t}$, is absorbed by the 2D electrons. Our analysis shows that the entire flux of axion-induced microwave radiation is fully absorbed in the quantum Hall state. Through out the paper, the temperature is of the order of 1010mK so that chemical potential is nearly equal to the Fermi energy $E_{f}$. We use physical units, Boltzmann constant $k_{B}=1$, light velocity $c=1$ and $\hbar=1$.
II integer quantum hall effect
[…]
$\Phi(\vec{x},n,k)=C_{n}H_{n}\Big(\frac{x-l_{B}^{2}k}{l_{B}}\Big)\exp\Big(-\frac{(x-l_{B}^{2}k)^{2}}{2l_{B}^{2}}\Big)\exp(-iky)$ (4)
with normalization constant $C_{n}=(2^{n}n!\sqrt{\pi}l_{B}L_{y})^{-1/2}$, where $L_{y}$ is the length of the Hall bar in the direction $y$. $l_{B}=\sqrt{1/eB}$ is magnetic length; $l_{B}\simeq 8.1\times 10^{-7}\rm cm\sqrt{10^{5}\rm Gauss/B}$. Degenerate states in each Landau level are specified by $k$. ( The coordinate of the two dimensional space is $x$ and $y$ and $\vec{B}=(0,0,B)$.)
Including the spin contribution, each Landau level is split to two states with energies $E_{n\pm}=\omega_{c}(n+1/2)\pm g\mu_{B}B$; $\pm$ correspond to spin parallel and anti-parallel to the magnetic field $B$, respectively. Namely, Zeeman energy $\pm g\mu_{B}B$ with $g\simeq 0.44$ and Bohr magneton $\mu_{B}=e/2m_{e}$. Typically, it is of the order of $10^{-4}(B/10^{5}\rm Gauss)$eV. Obviously Zeeman energy is smaller than the cyclotron energy in semiconductor. The degeneracy of each Landau level with definite spin is given by $eB/2\pi$.
III localization of two dimensional electrons
In general, real samples contain a disorder potential $V$, which lifts the degeneracy of energy levels. It is well established that, in the presence of such disorder, all 2D electron states become localized localization . However, when a strong external magnetic field is applied, a portion of these electronic states remain delocalized. This partial delocalization is essential for the emergence of the quantum Hall effect.
Taking the disorder potential into account, the density of states ( DOS ) acquires a finite width. Numerical simulations show that the DOS, denoted $\rho(E)$, has a finite spread $\Delta E$ centered around energy levels $E_{n\pm}$, as schematically illustrated in Fig.(1). For example, the DOS may take the form
$\rho(E)\propto\exp\left(-\frac{(E-E_{n\pm})^{2}}{(\Delta E)^{2}}\right),$ (5)
with $E_{n\pm}\gg\Delta E$ andouemura . This form will be used in our later discussions. The disorder potential is assumed to contain equal amounts of attractive and repulsive contributions, resulting in a symmetric DOS profile centered at $E_{n\pm}$, as shown in the figure.
Figure 1: Density of state $\rho(E)$ in finite sample. There is a region in which states are extended over the sample, while states in other region are localized. Dotted line represents Fermi energy $E_{f}$
The width $\Delta E$ is assumed to be much smaller than the cyclotron energy $\omega_{c}=eB/m^{\ast}$ under a strong magnetic field $B$. This condition is satisfied when the disorder potential energy is significantly smaller than the cyclotron energy, i.e., $V\ll\omega_{c}\simeq 1.8\times 10^{-2}(B/10^{5}~\text{Gauss})~\text{eV}$.
Electrons occupy states with energies lower than the Fermi energy $E_{f}$, while states with energies higher than $E_{f}$ remain unoccupied. The system is characterized by the filling factor $\nu\equiv 2\pi\rho_{e}/eB$, where $\rho_{e}$ is the electron density. When $\nu<1$, only a portion of the lowest Landau level is filled. In contrast, when $\nu>1$, the lowest Landau level is fully occupied, and electrons begin to populate the second Landau level. Taking spin into account, all electrons still reside in the lowest Landau level as long as $\nu<2$.
In Fig.(1), we also indicate the energy range of extended states $E_{\beta}$, defined by $E_{c}-\delta\leq E_{\beta}\leq E_{c}+\delta$, where $E_{c}\equiv E_{n\pm}$. This region can be determined using a scaling law aoki1 ; aoki2 , which states that the localization (or coherence) length $\xi(E)$ diverges as
$\xi(E)\propto|E-E_{c}|^{-2.4}\quad\text{as }E\to E_{c}.$ (6)
Thus, the critical energies $E=E_{c}\pm\delta$ correspond to the condition where the coherence length $\xi(E)$ becomes comparable to the size of the sample. The coherence length $\xi(E)$ characterizes the spatial extent of a quantum state with energy $E$, and hence, within the interval $E_{c}-\delta,E_{c}+\delta]$, the states are effectively extended and contribute to electrical conduction. In infinitely large system, the interval vanishes, i.e. $\delta=0$. Only a state with energy $E_{c}$ is extended.
When the Fermi energy $E_{f}$ lies below $E_{c}-\delta$, electrons occupy only localized states and cannot carry current; the system is in an insulating phase. In this regime, the longitudinal conductivity $\sigma_{xx}=0$, and the transverse conductivity $\sigma_{xy}$, or equivalently, the Hall resistance $\rho_{xy}=1/\sigma_{xy}$, exhibits quantized plateaus as a function of magnetic field. These plateaus have quantized values, such as $\rho_{xy}=\Big(\frac{2\pi}{e^{2}}\Big)\Big(\frac{1}{n}\Big)\,\,\mbox{with}\,\,n\,\,\mbox{positive integer},$ arising from the topological nature of the quantum Hall effect topology1 ; topology2 .
On the other hand, when $E_{c}-\delta<E_{f}<E_{c}+\delta$, the system undergoes a plateau-to-plateau transition. In this regime, some electrons occupy extended states and can conduct electricity; the system behaves as a metal with $\sigma_{xx}\neq 0$. As a result, the conductivity $\sigma_{xy}$ and Hall resistance $\rho_{xy}$ change continuously with the magnetic field. It should be noted that the Fermi energy $E_{f}$ itself varies with the magnetic field $B$.
( In addition to insulating and metallic phases, quantum Hall systems can also exhibit a superconducting-like phase joseph1 ; joseph2 ; joseph3 ; iwazaki when two parallel quantum Hall layers are placed in close proximity. Josephson-like effects between such bilayer systems have indeed been experimentally observed joseph3 . )
In this paper, we focus solely on transitions between localized states induced by axion-generated radiations. That is, the Fermi energy is taken to satisfy $E_{f}<E_{c}-\delta-m_{a}$, where $m_{a}$ is the axion mass. Under these conditions, the longitudinal conductivity $\sigma_{xx}$ remains zero even in the presence of axion-induced radiation. Consequently, the system remains in the insulating phase.
IV axion dark matter
In order to discuss the absorption of the axion induced radiation in quantum Hall states, we explain electromagnetic coupling with axionaxion1 ; axion2 ; axion3 and smallness of the coupling. The coupling $g_{a\gamma\gamma}$ between axion $a(t,\vec{x})$ and electromagnetic field is given by $L_{a\gamma\gamma}=g_{a\gamma\gamma}a(t,\vec{x})\vec{E}\cdot\vec{B}$ with electric $\vec{E}$ and magnetic $\vec{B}$ fields. The coupling $g_{a\gamma\gamma}$ is rewritten such that $g_{a\gamma\gamma}=g_{\gamma}\alpha/f_{a}\pi$, where fine structure constant $\alpha\simeq 1/137$ and axion decay constant $f_{a}$ satisfying the relation $m_{a}f_{a}\simeq 6\times 10^{-6}\rm eV\times 10^{12}$GeV in the QCD axion. The parameter $g_{\gamma}$ depends on the axion model, i.e. $g_{\gamma}(\rm DFSZ)\simeq 0.37$ for DFSZ modeldfsz ; dfsz1 and $g_{\gamma}(\rm KSVZ)\simeq-0.96$ for KSVZ modelksvz ; ksvz1 . Hereafter we only consider QCD axion model.
The coupling is much small for the axion dark matter which is described by $a(t,\vec{x})\simeq a_{0}\cos(m_{a}t)$. ( The momentum $k$ of the axion dark matter is of the order of $10^{-3}m_{a}$ near the Earth. We may neglect the momentum. ) Supposing that the local energy density $\rho_{d}$ of the dark matter is composed of the axion, $\rho_{d}=m_{a}^{2}\overline{a(t,\vec{x})^{2}}=m_{a}^{2}a_{0}^{2}/2\sim 0.3\rm GeV/cm^{3}$ ( $\overline{Q}$ denotes time average of the quantity $Q$. ) Then, we find the coupling $g_{a\gamma\gamma}a(t,\vec{x})\sim 10^{-21}$ is very small and it is independent of the axion mass in the QCD axion model.
Owing to the coupling, we have oscillating electric field $E_{a}=g_{a\gamma\gamma}a(t)B$ in the presence of the axion under the magnetic field $B$. The electric field makes electrons in metal oscillate so that oscillating electric current in the metal produce electromagnetic radiations. This is a mechanism of the production of electromagnetic radiations in resonant cavity under strong magnetic field. The radiation is represented by the gauge field $\vec{A}_{a}=\vec{A}_{0}\exp(-i\omega t+i\vec{p}\cdot\vec{x})\simeq\vec{A}_{0}\exp(-im_{a}t)$ whose amplitude $|\vec{A}_{0}|\sim g_{a\gamma\gamma}a_{0}B/m_{a}$. When we appropriately tune the size of the cavity, the radiations are amplifiedsikivie ; iwazaki01 with the resonance. Because we have cosmological constraintsWil ; Wil1 ; Wil2 $m_{a}=10^{-6}\rm eV\sim 10^{-3}eV$ in QCD axion model, the electromagnetic radiations are microwaves. Hereafter we address the microwaves in the discussion.
To detect the amplified radiation, we utilize a quantum Hall system placed inside a resonant cavity. When a semiconductor sample exhibiting the quantum Hall effect absorbs axion induced microwaves amplified by resonance, its temperature increases. We can determine the presence of the axion and its mass with the detection of the temperature increase. This provides a new method for axion detection.
In the following, we calculate the absorption energy from the axion induced microwaves and estimate the resulting temperature increase in a sample exhibiting the quantum Hall effect.
V absorption of axion microwaves in quantum Hall state
To make our discussion more concrete, we consider a resonant cavity with parallel slabs, as shown in Fig.(2), following our previous work iwazaki01 . Let the distance between the two parallel, facing slabs be $l$, and assume that an external magnetic field $\vec{B}_{t}=(0,0,B_{t})$ is applied parallel to the slabs. By adjusting the distance $l$, the resonance condition can be satisfied.
Microwaves with electric field $\vec{E}_{a}=(0,0,E_{a})$ and magnetic field $\vec{B}_{a}=(B_{a},0,0)$ are generated in the presence of the external field $\vec{B}_{t}$. The resulting energy flux is $\vec{E}_{a}\times\vec{B}_{a}=(0,E_{a}B_{a},0)$, indicating propagation in the $x$-direction.
It should be noted that the fields $E_{a}$ and $B_{a}$ vary with the spatial coordinate $y$, assuming the slabs are located at $y=0$ and $y=l$. The spatial dependence at non-resonance is given by $E_{a}=-E_{a0}\cos(m_{a}y+\delta)/\cos\delta$ and $B_{a}=E_{a0}\sin(m_{a}y+\delta)/\cos\delta$, where $\tan\delta=\sin(m_{a}l)/(1+\cos(m_{a}l))$ and $E_{a0}=g_{a\gamma\gamma}a_{0}B_{t}$.
At resonance, $l=\pi/m_{a}$, the fields become $E_{a}=\sqrt{2}E_{a0}\sin(m_{a}y)/(m_{a}\delta_{e})$ and $B_{a}=-\sqrt{2}E_{a0}\cos(m_{a}y)/(m_{a}\delta_{e})$, where $\delta_{e}=\sqrt{2/(m_{a}\sigma)}$ is the electromagnetic penetration depth in the slabs, and $\sigma$ is their electrical conductivity. Note that at resonance, the fields are significantly amplified because $m_{a}\delta_{e}\ll 1$. Here we have neglected time dependence of these fields because the dependence has no important play in the discussion below.
Figure 2: Slabs parallel to each other separated with distance $l$. Sample tilted with $\theta=\pi/6$ to $y$ axis and parallel to $x$ axis.
To facilitate absorption of the microwave, the two-dimensional electron layer should be slightly tilted rather than aligned perpendicular to the magnetic field $\vec{B}_{t}$. This tilt ensures that the electric field of the axion-induced microwave has a component parallel to the layer. Only the component is absorbed by 2D electrons. In our setup, the surface of the sample is aligned along the $x$-direction, as shown in Fig.(2). We tentatively take the angle $\theta$ between the $y$-axis and the 2D electron layer to be $\pi/6$.
The applied magnetic field $\vec{B}_{t}$ thus has a perpendicular component $B=B_{t}\cos(\theta=\pi/6)\approx 0.86B_{t}$, and a parallel component $B_{p}=B_{t}\sin(\pi/6)=0.5B_{t}$. Even though $\vec{B}_{t}$ is not perpendicular to the 2D layer, the quantum Hall effect is governed by the perpendicular component $B$.
Before discussing microwave absorption by 2D electrons, we examine how the electromagnetic wave is refracted inside the semiconductor due to its electric permittivity. In GaAs (used in our analysis), the relative permittivity is approximately $\epsilon\simeq 13$, while $\epsilon_{0}=1$ in vacuum.
For an incident angle $\theta_{i}=\pi/2-\theta=\pi/3$, the refraction angle $\theta_{r}$ satisfies $\sqrt{\epsilon}\sin\theta_{r}=\sin\theta_{i}$, yielding $\sin\theta_{r}\approx 0.24$.
The strengths of the electric and magnetic fields inside the semiconductor are given by
$E^{\prime}_{a}=\frac{2\cos\theta_{i}}{\sqrt{\epsilon}\cos\theta_{i}+\cos\theta_{r}}E_{a}\approx 0.36E_{a},\quad B^{\prime}_{a}=\sqrt{\epsilon}E^{\prime}_{a}\approx 1.3E_{a}.$ (7)
The component of the electric field parallel to the 2D electron plane is $E^{\prime}_{p}=E^{\prime}_{a}\cos\theta_{r}\approx 0.97E^{\prime}_{a}$. The magnetic field $\vec{B}^{\prime}_{a}$ inside the semiconductor is fully parallel to the 2D layer. It is important to emphasize that the microwave field $E_{a}\propto g_{a\gamma\gamma}a_{0}B_{t}$ is the value generated in vacuum, not within the semiconductor.
The component of the energy flux perpendicular to the 2D electrons is $E^{\prime}_{p}B^{\prime}_{a}\approx 0.46E_{a}^{2},$ which is the portion absorbed by the 2D electron system in the quantum Hall state.
For reference, the relevant parameters are summarized as follows
$E^{\prime}_{p}\approx 0.35E_{a},\quad B_{a}^{\prime}\approx 1.3E_{a},\quad B\approx 0.86B_{t},\quad E^{\prime}_{p}B^{\prime}_{a}\approx 0.46E_{a}^{2}.$ (8)
The cyclotron frequency is $\omega_{c}=eB/m^{*}$, the magnetic length is $l_{B}=\sqrt{1/eB}$, while the Zeeman energy is $g\mu_{B}B_{t}$.
Next, we analyze the absorption of axion-induced microwaves by the 2D electrons in the quantum Hall state. A state in Landau level $n$ is described by the wave function $\Phi_{\alpha}(x,y)=e^{-iE_{\alpha}t}\int dk\,f_{\alpha}(k)\Phi(\vec{x},n,k),$ where $f_{\alpha}(k)$ is chosen such that $\Phi_{\alpha}$ is an eigenstate of the electron Hamiltonian including the disorder potential $V$. The normalization condition is $1=\int dx\,dy\,\overline{\Phi_{\alpha}(x,y)}\Phi_{\alpha}(x,y)=\int dk\,\overline{f_{\alpha}(k)}f_{\alpha}(k).$
We neglect mixing between different Landau levels and between different spin states.
Since the extension of the semiconductor sample is much smaller than $1/m_{a}\sim 2\text{ cm}(10^{-5}\text{ eV}/m_{a})$, we can ignore the $y$-dependent terms in $E_{a}$ and $B_{a}$, and approximate $E_{a}\approx E_{a0}=g_{a\gamma\gamma}a_{0}B_{t}$. This simplification does not affect the final results.
The absorption of the microwaves $\vec{A}^{\prime}_{a}\sim\vec{A}^{\prime}_{0}\exp(-im_{a}t)$ with $|\vec{A}^{\prime}_{0}|=E^{\prime}_{p}/m_{a}\simeq 0.35g_{a\gamma\gamma}a_{0}B_{t}/m_{a}$, is described by the interaction Hamiltonian, $H_{a}=\frac{-ie\vec{A}^{\prime}_{a}\cdot\vec{P}}{m^{\ast}}$ with $\vec{P}=-i\vec{\partial}+e\vec{A}_{ex}$ and $\vec{B}=\vec{\partial}\times\vec{A}_{ex}$.
The matrix element between states $\alpha$ and $\beta$ is $\langle\beta|H_{a}|\alpha\rangle=i(E_{\beta}-E_{\alpha})e\vec{A}^{\prime}_{a}\cdot\vec{L}_{\alpha\beta},\quad\vec{L}_{\alpha\beta}\equiv\langle\beta|\vec{x}|\alpha\rangle.$
The transition rate $\tau_{t}^{-1}$ of an electron per unit time is
$\tau_{t}^{-1}=2\pi S\int dE_{\beta}\,\rho(E_{\beta})|\langle\beta|H_{a}|\alpha\rangle|^{2}\delta(E_{\alpha}+m_{a}-E_{\beta})\approx 2\pi S\rho(E_{\alpha}+m_{a})m_{a}^{2}(e\vec{A}^{\prime}_{a}\cdot\vec{L}_{\alpha\beta})^{2}.$ (9)
Assuming $|\vec{A}^{\prime}_{a}\cdot\vec{L}_{\alpha\beta}|^{2}\approx A_{0}^{\prime 2}A^{2}l_{B}^{2}$, we write
$\tau_{t}^{-1}\sim 2\pi S\rho(E_{\alpha}+m_{a})m_{a}^{2}e^{2}A_{0}^{\prime 2}A^{2}l_{B}^{2}.$ (10)
with surface area $S$ of the semiconductor sample. The density of states is modeled, for example, as
$\rho(E)=\rho_{0}\exp\left(-\left(\frac{E-E_{c}}{\Delta E}\right)^{2}\right),\quad\rho_{0}=\frac{eB}{2\pi}\cdot\frac{2}{\sqrt{\pi}\Delta E},$ (11)
with $E_{c}=\omega_{c}/2$, where $\int dE\rho(E)/2=eB/2\pi$ represents the number density of electrons in a Landau level.
The width $\Delta E$ is on the order of $\sim 10^{-3}$eV. It can be estimated using the approximate formula $\Delta E=\sqrt{\frac{2\omega_{c}}{\pi\tau_{r}}},$ where $\tau_{r}$ denotes the relaxation time of electrons. This relaxation time can be expressed in terms of the mobility $\mu$ of the semiconductor as $\tau_{r}=\mu m_{e}^{\ast}/e.$ Typically, $\tau_{r}\simeq 3.8$ps for $\mu=10^{5}\rm cm^{2}/Vs$. Substituting this into the expression for $\Delta E$, we find $\Delta E\simeq 1.4\times 10^{-3}\text{eV}\sqrt{\left(B/10^{5}\text{Gauss}\right)\left(10^{5}\text{cm}^{2}(\text{Vs})^{-1}/\mu\right)}.$
The maximal number of the electrons able to absorb axion induced microwave is
$\Delta N=S\int^{E_{f}}_{E_{f}-m_{a}}dE\rho(E)\simeq S\rho_{0}m_{a}\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right),$ (12)
leading to the maximal number of electrons absorbing microwaves per unit time
$\Delta N\tau_{t}^{-1}\sim 2\pi S^{2}\rho_{0}^{2}m_{a}^{3}e^{2}A_{0}^{\prime 2}A^{2}l_{B}^{2}\exp\left(-2\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right).$ (13)
We have used $m_{a}=10^{-5}$$\sim 10^{-4}eV$ $\ll E_{f}\sim E_{c}\sim 10^{-2}$eV.
We compared it to the number of axion-induced photons incoming per unit time, $0.46SE_{a}^{2}/m_{a}\approx 3.8Sm_{a}A_{0}^{\prime 2}.$ We find that full absorption occurs when
$\frac{\Delta N\tau_{t}^{-1}}{3.8Sm_{a}A_{0}^{\prime 2}}\simeq 6.7\times 10^{5}\Big(\frac{S}{1\rm cm^{2}}\Big)\Big(\frac{B}{10^{5}\rm Gauss}\Big)\Big(\frac{m_{a}}{10^{-5}\rm eV}\Big)^{2}\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)^{2}A^{2}\exp\Big(-2\Big(\frac{E_{f}-E_{c}}{\Delta E}\Big)^{2}\Big)>1.$ (14)
The condition can be satisfied unless the overlapping factor $A$ ( $\beta|x|\alpha>=Al_{B}$ ) is not too small, e.g. $A\ll 1$.
Namely, all photons $3.8Sm_{a}A_{0}^{\prime 2}$ incoming are absorbed within a time $\tau_{t}$, because there are sufficiently number of electrons $\Delta N$ which are able to absorb the photons within the time $\tau_{t}$. The condition in eq(14) holds even for external microwave irradiation with frequency $m_{a}$ and any power. Then, it apparently seems that incoming microwaves with any power are fully absorbed.
However, for full absorption of microwaves, the incoming microwave power, which is proportional to $E^{2}$ (where $E$ is the electric field), must be not too large. If the microwave power becomes high enough that the transition time $\tau_{t}\propto 1/E^{2}$ becomes shorter than the relaxation time $\tau_{r}$ of the 2D electrons, then the microwaves may not be fully absorbed.
Specifically, when $\tau_{t}>\tau_{r}$, the electrons that have been excited to higher-energy states can relax and thermalize within time $\tau_{r}$. This allows the absorption process to repeat: localized electrons with energy $E_{\alpha}<E_{f}$ can absorb microwaves and transition to localized states with $E_{\beta}>E_{f}$.
In contrast, when $\tau_{t}\ll\tau_{r}$, the excited electrons do not have time to thermalize, and thus do not return to their original localized states within time $\tau_{t}$. As a result, those states with energy $E_{\alpha}$ ( $E_{f}-m_{a}<E_{\alpha}<E_{f}$ ) remain vacant so that electrons with energies lower than $E_{f}-m_{a}$ should absorb the microwaves. But the number of such electrons with lower energies are much less than $\Delta N$. Eventually, there are no electrons available to absorb additional microwaves. ( Some of the incoming microwaves may be absorbed by excited electrons with energies above the Fermi energy $E_{f}$ which have already absorbed previous microwave photons. However, it remains unclear whether these excited electrons can fully absorb the incoming microwave energy. )
Therefore, in cases where high-power microwaves are incident on the 2D electron system, a portion of the radiation may not be absorbed and can transmit through the sample.
In the case of the axion induced microwave, we can show that they are fully absorbed because the transition time $\tau_{t}$ is much longer than the relation time even if the power of the microwave is amplified with resonance.
$\tau_{t}\sim 10^{-3}\text{ s}\left(\frac{1\text{ cm}^{2}}{S}\right)\left(\frac{m_{a}}{10^{-5}\text{ eV}}\right)\left(\frac{10^{5}\text{ Gauss}}{B}\right)^{2}\left(\frac{1}{A^{2}}\right)\left(\frac{0.3\,\text{GeV}/\text{cm}^{3}}{\rho_{d}}\right)\exp\left(\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right),$ (15)
which is much longer than relaxation time of the order of $\tau_{r}\sim 10^{-12}\text{ s}$, unless $A$ is much large $\gg 1$.
Hence, axion-induced microwaves are fully absorbed by the 2D electron gas in the quantum Hall regime. As a result, the sample’s temperature increases, albeit only slightly in the absence of resonance. The incoming power is tiny
$P_{ra}\sim 0.46SE_{a}^{2}\sim 10^{-30}\text{ W}\left(\frac{B}{10^{5}\text{ Gauss}}\right)^{2}\left(\frac{S}{1\text{cm}^{2}}\right)\left(\frac{\rho_{d}}{0.3\,\text{GeV}/\text{cm}^{3}}\right),$ (16)
but it can be amplified to $\sim 10^{-18}\text{ W}$ using a properly designed resonant cavity.
VI temperature increase of sample
We show that the absorption of the power $P_{ra}$, estimated above, slightly increases the temperature of the sample exhibiting the quantum Hall effect. However, this power is not large enough to produce a detectable temperature rise.
Now, we consider the effect of resonance. We demonstrate that the temperature increase becomes sufficiently large to be detectable when resonance occurs in the cavity, achieved by tuning an appropriate length scale.
In the case of two parallel slabs, ( see Fig.(2) ), resonance takes place when the separation between the slabs is $l=\frac{\pi}{m_{a}}\simeq 6.3\,\mbox{cm}\big(\frac{10^{-5}\,\rm{eV}}{m_{a}}\big)$.
When resonance occurs, the axion-induced microwave field in vacuum (with $B_{a}\sim E_{a}$) is amplified iwazaki01 as
$E_{a}\to\frac{E_{a}}{m_{a}\delta_{e}},\quad\text{with penetration depth}\quad\delta_{e}=\sqrt{\frac{2}{m_{a}\sigma}},$ (17)
where $\sigma$ is the electrical conductivity of the metal forming the cavity (the slabs). For example, for 6N copper at temperatures below 1 K, $\sigma\simeq 3.3\times 10^{7}\,\text{eV}$ (corresponding to $\sigma^{-1}=0.2\times 10^{-11}\,\Omega\cdot\text{m}$).
This leads to the enhancement factor
$\frac{1}{m_{a}\delta_{e}}\simeq 1.3\times 10^{6}\sqrt{\left(\frac{\sigma}{3.3\times 10^{7}\,\text{eV}}\right)\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)}.$ (18)
The electric field $E^{\prime}_{p}$ in the semiconductor, as well as the magnetic field $B^{\prime}_{a}$, is enhanced by the same factor $1/(m_{a}\delta_{e})$.
As a result, the amplified energy flux entering the sample is
$\frac{E^{\prime}_{p}B^{\prime}_{a}}{(m_{a}\delta_{e})^{2}}\simeq 0.46\left(\frac{E_{a}}{m_{a}\delta_{e}}\right)^{2},$ (19)
and it is absorbed in the quantum Hall state. The power absorbed by the sample with surface area $S$ is given by
$P_{ra}\sim 0.46\left(\frac{E_{a}}{m_{a}\delta_{e}}\right)^{2}S\sim 0.9\times 10^{-18}\,\text{W}\,g_{\gamma}^{2}\frac{S}{\text{cm}^{2}}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV/cm}^{3}}\right),$ (20)
with $E_{a}^{2}=(g_{a\gamma\gamma}a_{0}B_{t})^{2}\simeq 4.0\times 10^{-30}g_{\gamma}^{2}\,\text{W/cm}^{2}\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}$.
It should be contrasted with the power absorbed in conductor with large electrical conductivity. Electric field in the conductor is suppressed by factor $m_{a}\delta_{e}$iwazaki01 ; kishimoto ; iwazaki02 compared with the one in vacuum. Thus, the power is enhanced only by the factor $1/m_{a}\delta_{e}\sim 10^{6}$ when the resonance occurs.
The temperature of the sample increases due to this absorbed energy. To estimate the temperature rise, we need to know the heat capacities of both the semiconductor and the 2D electrons. According to Debye model, the heat capacity $C_{s}$ of a rectangular semiconductor with volume $Sd$ at low temperatures (below 100 mK) is
$C_{s}\simeq 1.94\times 10^{3}\,\text{J/g\textperiodcentered K}\left(\frac{T}{T_{D}}\right)^{3}\left(\frac{\text{density}}{M}\right)Sd,$ (21)
where $M$ is the molecular weight of the semiconductor and $T_{D}$ is the Debye temperature. For GaAs, $M=144$, density = $5.3\,\text{g/cm}^{3}$, and $T_{D}=360\,\text{K}$. (We assume that the semiconductor’s heat capacity is dominated by phonons.)
Numerically, this gives
$C_{s}\simeq 0.88\times 10^{7}\left(\frac{T}{20\,\text{mK}}\right)^{3}\left(\frac{S}{1\,\text{cm}^{2}}\right)\left(\frac{d}{0.1\,\mu\text{m}}\right).$ (22)
On the other hand, the heat capacity of 2D electrons in the low-temperature limit is
$C_{e}=\frac{dQ}{dT}\simeq\frac{\pi^{2}}{3}TS\rho(E_{f})=\frac{\pi^{2}}{3}TS\rho_{0}\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right),$ (23)
where $Q$ is the energy of the 2D electron system,
$Q=\int_{0}^{\infty}dE\,E\,\rho(E)\frac{1}{\exp\left(\frac{E-E_{f}}{T}\right)+1}.$ (24)
Numerically, this yields
$C_{e}\simeq 3.4\times 10^{8}\left(\frac{S}{1\,\text{cm}^{2}}\right)\left(\frac{T}{20\,\text{mK}}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{1/2}\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right).$ (25)
Thus, when $\exp\left(-(E_{f}-E_{c})^{2}/\Delta E^{2}\right)<10^{-2}$, the dominant contribution to the total heat capacity is from the semiconductor, i.e., $C_{s}+C_{e}\approx C_{s}$.
However, note the condition in Eq.(14), which requires $\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)A\exp\left(-(E_{f}-E_{c})^{2}/\Delta E^{2}\right)$ not to be too small. This condition must be satisfied for the axion-induced microwave to be completely absorbed.
Therefore, it is likely that choosing the Fermi energy $E_{f}$ such that
$10^{-3}\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)^{-1}A^{-1}<\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right)<10^{-2}$ (26)
which ensures both that $C_{s}>C_{e}$ and that the microwave is effectively absorbed. It is favorable to have smaller width $\Delta E<10^{-3}$eV, which is realized with larger mobility $\mu>10^{5}\rm cm^{2}/Vs$. The Fermi energy is approximately in the range $E_{f}\simeq E_{c}-(1.5\sim 2.6)\Delta E$ with $\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)A=1$, that is, in the tail of the density of states $\rho(E_{f})$.
In actual experiments, we control not the Fermi energy $E_{f}$ directly, but the external magnetic field $B_{t}$. To meet the conditions described above, we must choose the magnetic field such that the resulting Fermi energy falls within the required range.
For example, with $B_{t}=1.5\times 10^{5}\,\text{Gauss}$, we find $E_{c}=\omega_{c}/2\simeq 2.3\times 10^{-2}\,\text{eV}$, and the Zeeman energy is about $3.9\times 10^{-4}\,\text{eV}$. Since the Zeeman splitting is much smaller than the cyclotron energy $\omega_{c}$, the quantum Hall plateaus at $\nu=1$ and $\nu=2$ are nearly degenerate. We must appropriately choose the magnetic field so that $E_{f}$ is near the edge (i.e., $\nu<1$) of the $\nu=1\sim 2$ plateau in $\rho_{xy}$ or $\sigma_{xy}$.
In order to find appropriate magnetic field, we test the temperature response of the sample under weak external microwave irradiation with power on the order of $\sim 10^{-18}\,\text{W}$. By appropriately tuning the magnetic field $B_{t}$, full absorption arises, resulting in a detectable temperature rise. This preparatory step helps us optimize the strength of the magnetic field, in other words, Fermi energy $E_{f}$.
Now, when the sample absorbs the amplified microwave radiation, the rate of temperature increase is given by
$\dot{T}=\frac{P_{ra}}{C_{s}(T)}\simeq\frac{7.2\,\text{mK}}{\text{s}}\,g_{\gamma}^{2}\left(\frac{20\,\text{mK}}{T}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}/\text{cm}^{3}}\right)$ (27)
where we take the thickness $d=0.1\,\mu\text{m}$ and temperature $T=20\,\text{mK}$, with an external magnetic field $B_{t}=1.5\times 10^{5}\,\text{Gauss}$, applied parallel to the sample layers. ( Even when smaller magnetic field $B_{t}=6\times 10^{4}$Gauss, $\dot{T}\simeq 0.032$mK/ms with lower temperature T=10T=10mK. )
The thermal energy generated from microwave absorption dissipates into the thermal bath. In the actual experiment, we measure the temperature increase in the presence of this thermal dissipation.
The temperature of the sample, $T+\Delta T$ (with $T=20\,\text{mK}$), evolves according to the differential equation
$C_{s}\frac{d\Delta T}{dt}=-G\Delta T+P_{ra}$ (28)
with initial condition $\Delta T(t=0)=0$, where $G$ denotes the thermal conductance. Since $C_{s}\propto T^{3}$ and $G\propto T^{3}$, the stationary solution is
$\Delta T(t=\infty)=\frac{P_{ra}\tau}{C_{s}(\Delta T(t=\infty)+20\,\text{mK})}\simeq\frac{P_{ra}\tau}{C_{s}(20\,\text{mK})}\,\,\,\text{for}\,\,\,\Delta T\ll 20\text{mK}$ (29)
where $\tau\equiv C_{s}/G$ is the time constant such that without microwave power ($P_{ra}=0$), the excess temperature decays as $\Delta T\propto\exp(-t/\tau)$ as $t\to\infty$.
Thus, the final temperature increase due to axion-induced microwave absorption is approximately
$\Delta T(t=\infty)\simeq$ $0.72$ $\,\text{mK}\left(\frac{\tau}{100\,\text{ms}}\right)\,g_{\gamma}^{2}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)$ (30)
$\times$ $\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$
The formula is applicable only for QCD axion model in which the coupling constant $g_{a\gamma\gamma}$ satisfies $g_{a\gamma\gamma}\simeq 39g_{\gamma}(10^{-14}\text{GeV}^{-1})(m_{a}/10^{-5}\text{eV})$. Using model independent coupling constant $g_{a\gamma\gamma}$, we represent the temperature increase in the following,
$\Delta T(t=\infty)\simeq$ $5$ $\text{mK}\left(\frac{\tau}{100\,\text{ms}}\right)\,\Big(\frac{g_{a\gamma\gamma}}{10^{-14}\text{GeV}^{-1}}\Big)^{2}\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)$ (31)
$\times$ $\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$
We show in Fig.(3) the excluded region of $g_{a\gamma\gamma}$ and $m_{a}$ indicated by dashed line when we do not observe $\Delta T$ equal to or larger than $0.5\text{mK}(\tau/10\text{ms})$ in the case of time constant $\tau=10$ms. If the time constant $\tau$ is less than 1010ms, for instance, 11ms, the excluded region is above the same dashed line indicated with $\Delta T=0.05\text{mK}(\tau/1\text{ms})$ as the line in the figure when we do not observe $\Delta T\geq 0.05$mK. In the figure, we have roughly depicted the excluded region by CAPPCAPP1 ; CAPP2 ; CAPP3 . We also show the lines associated with the models of QCD axion.
Figure 3: When we do not observe $\Delta T$ larger than $0.5\text{mK}(\tau/10\text{ms})$ with time constant $\tau=10$ms, the value of $g_{a\gamma\gamma}$ above the dashed line denoted with $0.5\text{mK}(\tau/10\text{ms})$ is excluded. We roughly show excluded region obtained by CAPP.
Therefore, we find that a large time constant $\tau>10\,\text{ms}$ is preferable to observe a measurable temperature increase $\Delta T$. To achieve such a large time constant, it is needed to cover the semiconductor sample with a transparent, thin sheet that has both lower thermal conductivity and lower heat capacity than those of the sample semiconductor. Additionally, minimizing the heat capacity $C_{s}$ by using a very thin sample (e.g., $d\leq 0.1\,\mu\text{m}$) is advantageous. Importantly, the result $\Delta T$ does not depend on the surface area $S$ of the sample since both $P_{ra}\propto S$ and $C_{s}\propto S$; $\Delta T\propto P_{ra}/C_{s}$.
To identify the resonance condition, we finely tune the cavity length $l$. Since the axion energy $m_{a}+\frac{1}{2}m_{a}v^{2}$ (with $v\sim 10^{-3}$) is distributed around $m_{a}$ with a relative width of $\sim 10^{-6}$, the cavity length must be scanned in steps of $\delta l=10^{-6}(\pi/m_{a})\simeq 6.3\times 10^{-6}\,\text{cm}\left(10^{-5}\,\text{eV}/m_{a}\right)$.
If we measure the sample temperature for 1 second at each step, scanning the axion mass range $(0.9\sim 1.1)\times 10^{-5}\,\text{eV}$ takes approximately 50 hours. While the measurement time per step may vary depending on the detection method, this approach provides a highly efficient means of searching for axions.
Our idea for the detection of dark matter axion is that we measure the increase of the temperature caused by the absorption of much weak microwaves generated by the axion. So it is favorable to use materials with much small heat capacity. In the paper we have proposed two dimensional electrons forming quantum Hall states as such a material. But the idea is also applicable for materials which are able to absorb the microwaves and have sufficiently small heat capacities for the temperature to increase. To find such a material is a subject in near future.
VII conclusion
We have proposed a novel method for axion detection that leverages the quantum Hall effect within an axion haloscope. By precisely tuning the haloscope’s dimensions, axion-induced microwave radiation can be resonantly amplified. This amplified radiation is absorbed by a semiconductor sample exhibiting the quantum Hall effect, leading to a measurable increase in temperature—a signal that would otherwise be undetectable without amplification.
[…]
Although the method for measuring the temperature rise is beyond the scope of this work, the predicted increase offers a direct and sensitive means for detecting the axion and determining its mass.
The author expresses thanks to A. Sawada for useful comments.

### v3

@@
 By monitoring this temperature increase, the mass $m_{a}$ of the axion can be inferred.
-As an example, consider a GaAs sample with small thickness $d=0.1\,\mu\mathrm{m}$ and its heat capacity $C_{s}$ at temperature $T=20\,\mathrm{mK}$.
-According to the QCD axion model, because the energy flux of the incoming radiation is $P_{ra}\sim 10^{-18}g_{\gamma}^{2}\mbox{W cm}^{-2}(10^{-5}\mbox{eV}/m_{a})(B/1.5\times 10^{5}\mbox{Gauss})^{2}(\rho_{d}/0.3\rm GeVcm^{-3})$ at the resonance with $g_{\gamma}(\rm KSVZ)=-0.96$ and $g_{\gamma}(\rm DFSZ)=0.37$, the temperature increase $(P_{ra}/C_{s})\times\tau$ is approximately $\simeq 0.25(\mbox{mK})(\tau/10\mbox{ms})\,g_{\gamma}^{2}\,(20\mbox{mK}/T)^{3}(10^{-5}\mbox{eV}/m_{a})(B/1.5\times 10^{5}\mbox{Gauss})^{2}(0.1\mu\mbox{m}/d)$, where $\tau$ is the time constant associated with the heat dissipation into thermal bath.
+As an example, consider a GaAs sample with small thickness $d=1\,\mu\mathrm{m}$ and its heat capacity $C_{s}$ at temperature $T=20\,\mathrm{mK}$.
+According to the QCD axion model, because the energy flux of the incoming radiation is $P_{ra}\sim 10^{-18}g_{\gamma}^{2}\mbox{W cm}^{-2}(10^{-5}\mbox{eV}/m_{a})(B/1.5\times 10^{5}\mbox{Gauss})^{2}(\rho_{d}/0.3\rm GeVcm^{-3})$ at the resonance with $g_{\gamma}(\rm KSVZ)=-0.96$ and $g_{\gamma}(\rm DFSZ)=0.37$, the temperature increase $(P_{ra}/C_{s})\times t_{ob}$ is approximately $\simeq 0.72(\mbox{mK})(t_{ob}/1\text{s})\,g_{\gamma}^{2}\,(20\mbox{mK}/T)^{3}(10^{-5}\mbox{eV}/m_{a})(B/1.5\times 10^{5}\mbox{Gauss})^{2}(1\mu\text{m}/d)$, where $t_{ob}$ is the observation time.
+It must be smaller than a time constant $\tau$ associated with the heat dissipation into thermal bath.
+Such a large time constant should be realized by making thermal conductance much small.
 I introduction
@@
 This temperature rise serves as an observable signature of axion presence and allows for the determination of the axion mass.
+For larger axion masses $m_{a}$, the cavity volume scales as $m_{a}^{-3}$ and therefore becomes smaller, leading to a corresponding reduction in the axion-induced power, which also scales as $m_{a}^{-3}$.
+As a result, direct power measurements in conventional axion haloscopes become increasingly challenging at higher axion masses.
+At resonance, the energy density of the amplified radiation scales as $m_{a}^{-1}$ and thus also decreases with increasing axion mass, but much more gradually than the axion-induced power.
+Consequently, the local energy absorbed by a small semiconductor sample is proportional to $m_{a}^{-1}$, and the resulting temperature rise rate follows the same scaling.
+This weaker dependence on $m_{a}$ makes temperature-based measurements a more effective approach for axion detection at higher masses.
+We note that the surface area of the sample is of the order of $(0.1\text{cm})^{2}$ or less.
+Only a fraction of induced radiations by the axion in the cavity is absorbed.
+Putting such a small sample in the resonant cavity, we can determine the presence of the axion and its mass with the detection of the temperature increase.
+This is an advantage over previous axion haloscopes.
 ( In previous papersiwaqhe1 ; iwaqhe2 ; iwaqhe3 we have pointed out that among the experiments conducted so far on the quantum Hall effect, there exist phenomena that appear to be influenced by dark matter axions.
@@
 By appropriately tuning the Fermi energy, the contribution from the 2D electrons $C_{e}$ can be made negligible compared to $_{s}; CC$$\simeq C_{s}$.
+Or simply by taking the volume, e.g.
+thickness of the semiconductor sample large, heat capacity $C_{s}$ can be larger than $C_{e}$.
 At low temperatures (e.g., 1010mK), the semiconductor heat capacity $C_{s}$ is dominated by phonons and follows the relation $C_{s}\propto T^{3}$, making the temperature increase highly sensitive to the temperature $T$.
@@
 As a concrete example, for a GaAs sample involving the quantum Hall state, the temperature increase rate at resonance is approximately
-$\frac{P_{ra}}{C}\simeq\frac{0.025\,\mathrm{mK}}{\mathrm{ms}}\times g_{\gamma}^{2}\left(\frac{20\,\mathrm{mK}}{T}\right)^{3}\Big(\frac{10^{-5}\mbox{eV}}{m_{a}}\Big)\left(\frac{B_{t}}{1.5\times 10^{5}\,\mathrm{Gauss}}\right)^{2}\left(\frac{0.1\,\mu\mathrm{m}}{d}\right)\Big(\frac{\rho_{d}}{0.3\rm GeVcm^{-3}}\Big),$ (1)
+$\frac{P_{ra}}{C}\simeq\frac{0.72\,\mathrm{mK}}{\mathrm{s}}\times g_{\gamma}^{2}\left(\frac{20\,\mathrm{mK}}{T}\right)^{3}\Big(\frac{10^{-5}\mbox{eV}}{m_{a}}\Big)\left(\frac{B_{t}}{1.5\times 10^{5}\,\mathrm{Gauss}}\right)^{2}\left(\frac{1\,\mu\mathrm{m}}{d}\right)\Big(\frac{\rho_{d}}{0.3\rm GeVcm^{-3}}\Big),$ (1)
 with the density $\rho_{d}$ of the dark matter axion, where $d$ is the sample thickness.
@@
 The form of semiconductor sample is rectangular with surface area $S$ and thickness $d$.
-Finally, because the absorbed energy dissipates into the thermal bath, the actual observed temperature increase $\Delta T$ depends on the time constant $\tau$ of the dissipation.
-For example, with the time constant $\tau=10$ms, we have
-$\Delta T\simeq 0.25\,\mathrm{mK}\times\left(\frac{\tau}{10\,\mathrm{ms}}\right)g_{\gamma}^{2}\left(\frac{20\,\mathrm{mK}}{T}\right)^{3}\Big(\frac{10^{-5}\mbox{eV}}{m_{a}}\Big)\Big(\frac{B_{t}}{1.5\times 10^{5}\mbox{Gauss}}\Big)^{2}\left(\frac{0.1\,\mu\mathrm{m}}{d}\right)\Big(\frac{\rho_{d}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
+It apparently seems that the temperature is larger as the observation time $t_{ob}$ is longer, $\Delta T=P_{ra}/C\times t_{ob}$.
+But, the absorbed energy dissipates into the thermal bath and the actual observed temperature increase $\Delta T$ depends on the time constant $\tau$ of the dissipation.
+However, the energy dissipation is very slow because the semiconductor sample is connected to thermal bath with small copper lines for measurement of electric current and small pedestal supposing the sample in dilution refrigerator.
+The surface area in contact with the thermal bath is much smaller than the one of the sample.
+Thus, the actual temperature increase is determined by the the observation time $t_{ob}$ as far as $t_{ob}<\tau$.
+As we discuss later, we may take $t_{ob}=1$ second and obtain for QCD axion,
+$\Delta T\simeq 0.72\,\mathrm{mK}\times\left(\frac{t_{ob}}{1\,\mathrm{s}}\right)g_{\gamma}^{2}\left(\frac{20\,\mathrm{mK}}{T}\right)^{3}\Big(\frac{10^{-5}\mbox{eV}}{m_{a}}\Big)\Big(\frac{B_{t}}{1.5\times 10^{5}\mbox{Gauss}}\Big)^{2}\left(\frac{1\,\mu\mathrm{m}}{d}\right)\Big(\frac{\rho_{d}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
 or for general coupling constant $g_{a\gamma\gamma}$,
-$\Delta T\simeq 0.5\text{mK}\left(\frac{\tau}{10\,\text{ms}}\right)\,\Big(\frac{g_{a\gamma\gamma}}{10^{-14}\text{GeV}^{-1}}\Big)^{2}\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)^{3}\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$ (3)
-It is favorable to cover the semiconductor sample with a transparent thin sheet possessing much small heat capacity and much small heat conductance for $\tau$ to be larger than 1010ms.
+$\Delta T\simeq 5\text{mK}\left(\frac{t_{ob}}{1\,\text{s}}\right)\,\Big(\frac{g_{a\gamma\gamma}}{10^{-14}\text{GeV}^{-1}}\Big)^{2}\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)^{3}\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{1\,\mu\text{m}}{d}\right)\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$ (3)
+It would be not difficult to fabricate the sample with small copper lines and small pedestal for time constant $\tau$ of the energy dissipation to be sufficiently longer than the observation time, i.e.
+11\,s.
 We begin by reviewing the fundamental properties of the quantum Hall state in 2D electron systems.
 We then demonstrate how microwave radiation, generated by axions in the presence of an external magnetic field $B_{t}$, is absorbed by the 2D electrons.
-Our analysis shows that the entire flux of axion-induced microwave radiation is fully absorbed in the quantum Hall state.
+Our analysis shows that the entire radiation flux entering the sample is fully absorbed in the quantum Hall state.
 Through out the paper, the temperature is of the order of 1010mK so that chemical potential is nearly equal to the Fermi energy $E_{f}$.
@@
 $g_{\gamma}(\rm DFSZ)\simeq 0.37$ for DFSZ modeldfsz ; dfsz1 and $g_{\gamma}(\rm KSVZ)\simeq-0.96$ for KSVZ modelksvz ; ksvz1 .
-Hereafter we only consider QCD axion model.
+Hereafter we mainly consider QCD axion model, although our results are also presented with general $g_{a\gamma\gamma}$ independent of $m_{a}$.
 The coupling is much small for the axion dark matter which is described by $a(t,\vec{x})\simeq a_{0}\cos(m_{a}t)$.
@@
 When a semiconductor sample exhibiting the quantum Hall effect absorbs axion induced microwaves amplified by resonance, its temperature increases.
-We can determine the presence of the axion and its mass with the detection of the temperature increase.
+The surface area of the sample is of the order of $(0.1\text{cm})^{2}$ or less so that only a fraction of induced radiations by the axion in the cavity is absorbed.
+Putting such a small sample in the resonant cavity, we can determine the presence of the axion and its mass with the detection of the temperature increase.
 This provides a new method for axion detection.
@@
 By adjusting the distance $l$, the resonance condition can be satisfied.
+We put a semiconductor sample involving two dimensional electrons in quantum Hall state.
+The sample in the cavity is very small compared with the length $l$.
+The surface area is of the order of $(0.1\text{cm})^{2}$ or less.
 Microwaves with electric field $\vec{E}_{a}=(0,0,E_{a})$ and magnetic field $\vec{B}_{a}=(B_{a},0,0)$ are generated in the presence of the external field $\vec{B}_{t}$.
@@
 We compared it to the number of axion-induced photons incoming per unit time, $0.46SE_{a}^{2}/m_{a}\approx 3.8Sm_{a}A_{0}^{\prime 2}.$ We find that full absorption occurs when
-$\frac{\Delta N\tau_{t}^{-1}}{3.8Sm_{a}A_{0}^{\prime 2}}\simeq 6.7\times 10^{5}\Big(\frac{S}{1\rm cm^{2}}\Big)\Big(\frac{B}{10^{5}\rm Gauss}\Big)\Big(\frac{m_{a}}{10^{-5}\rm eV}\Big)^{2}\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)^{2}A^{2}\exp\Big(-2\Big(\frac{E_{f}-E_{c}}{\Delta E}\Big)^{2}\Big)>1.$ (14)
+$\frac{\Delta N\tau_{t}^{-1}}{3.8Sm_{a}A_{0}^{\prime 2}}\simeq 6.7\times 10^{3}\Big(\frac{S}{\rm(0.1cm)^{2}}\Big)\Big(\frac{B}{10^{5}\rm Gauss}\Big)\Big(\frac{m_{a}}{10^{-5}\rm eV}\Big)^{2}\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)^{2}A^{2}\exp\Big(-2\Big(\frac{E_{f}-E_{c}}{\Delta E}\Big)^{2}\Big)>1.$ (14)
 The condition can be satisfied unless the overlapping factor $A$ ( $\beta|x|\alpha>=Al_{B}$ ) is not too small, e.g. $A\ll 1$.
-Namely, all photons $3.8Sm_{a}A_{0}^{\prime 2}$ incoming are absorbed within a time $\tau_{t}$, because there are sufficiently number of electrons $\Delta N$ which are able to absorb the photons within the time $\tau_{t}$.
+Namely, all photons $3.8Sm_{a}A_{0}^{\prime 2}$ incoming to the semiconductor sample are absorbed within a time $\tau_{t}$, because there are sufficiently number of electrons $\Delta N$ which are able to absorb the photons within the time $\tau_{t}$.
 The condition in eq(14) holds even for external microwave irradiation with frequency $m_{a}$ and any power.
@@
 In the case of the axion induced microwave, we can show that they are fully absorbed because the transition time $\tau_{t}$ is much longer than the relation time even if the power of the microwave is amplified with resonance.
-$\tau_{t}\sim 10^{-3}\text{ s}\left(\frac{1\text{ cm}^{2}}{S}\right)\left(\frac{m_{a}}{10^{-5}\text{ eV}}\right)\left(\frac{10^{5}\text{ Gauss}}{B}\right)^{2}\left(\frac{1}{A^{2}}\right)\left(\frac{0.3\,\text{GeV}/\text{cm}^{3}}{\rho_{d}}\right)\exp\left(\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right),$ (15)
+$\tau_{t}\sim 10^{-1}\text{ s}\left(\frac{\text{(0.1cm)}^{2}}{S}\right)\left(\frac{m_{a}}{10^{-5}\text{ eV}}\right)\left(\frac{10^{5}\text{ Gauss}}{B}\right)^{2}\left(\frac{1}{A^{2}}\right)\left(\frac{0.3\,\text{GeV}/\text{cm}^{3}}{\rho_{d}}\right)\exp\left(\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right),$ (15)
 which is much longer than relaxation time of the order of $\tau_{r}\sim 10^{-12}\text{ s}$, unless $A$ is much large $\gg 1$.
@@
 The incoming power is tiny
-$P_{ra}\sim 0.46SE_{a}^{2}\sim 10^{-30}\text{ W}\left(\frac{B}{10^{5}\text{ Gauss}}\right)^{2}\left(\frac{S}{1\text{cm}^{2}}\right)\left(\frac{\rho_{d}}{0.3\,\text{GeV}/\text{cm}^{3}}\right),$ (16)
-but it can be amplified to $\sim 10^{-18}\text{ W}$ using a properly designed resonant cavity.
+$P_{ra}\sim 0.46SE_{a}^{2}\sim 10^{-32}\text{ W}\left(\frac{B}{10^{5}\text{ Gauss}}\right)^{2}\left(\frac{S}{\text{(0.1cm)}^{2}}\right)\left(\frac{\rho_{d}}{0.3\,\text{GeV}/\text{cm}^{3}}\right),$ (16)
+but it can be amplified to $\sim 10^{-20}\text{ W}$ using a properly designed resonant cavity.
 VI temperature increase of sample
@@
 and it is absorbed in the quantum Hall state. The power absorbed by the sample with surface area $S$ is given by
-$P_{ra}\sim 0.46\left(\frac{E_{a}}{m_{a}\delta_{e}}\right)^{2}S\sim 0.9\times 10^{-18}\,\text{W}\,g_{\gamma}^{2}\frac{S}{\text{cm}^{2}}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV/cm}^{3}}\right),$ (20)
+$P_{ra}\sim 0.46\left(\frac{E_{a}}{m_{a}\delta_{e}}\right)^{2}S\sim 0.9\times 10^{-20}\,\text{W}\,g_{\gamma}^{2}\Big(\frac{S}{\text{(0.1cm)}^{2}}\Big)\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV/cm}^{3}}\right),$ (20)
 with $E_{a}^{2}=(g_{a\gamma\gamma}a_{0}B_{t})^{2}\simeq 4.0\times 10^{-30}g_{\gamma}^{2}\,\text{W/cm}^{2}\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}$.
@@
 Numerically, this gives
-$C_{s}\simeq 0.88\times 10^{7}\left(\frac{T}{20\,\text{mK}}\right)^{3}\left(\frac{S}{1\,\text{cm}^{2}}\right)\left(\frac{d}{0.1\,\mu\text{m}}\right).$ (22)
+$C_{s}\simeq 0.88\times 10^{6}\left(\frac{T}{20\,\text{mK}}\right)^{3}\left(\frac{S}{\text{(0.1cm)}^{2}}\right)\left(\frac{d}{1\,\mu\text{m}}\right).$ (22)
 On the other hand, the heat capacity of 2D electrons in the low-temperature limit is
@@
 Numerically, this yields
-$C_{e}\simeq 3.4\times 10^{8}\left(\frac{S}{1\,\text{cm}^{2}}\right)\left(\frac{T}{20\,\text{mK}}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{1/2}\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right).$ (25)
-Thus, when $\exp\left(-(E_{f}-E_{c})^{2}/\Delta E^{2}\right)<10^{-2}$, the dominant contribution to the total heat capacity is from the semiconductor, i.e., $C_{s}+C_{e}\approx C_{s}$.
+$C_{e}\simeq 3.4\times 10^{6}\left(\frac{S}{\text{(0.1cm)}^{2}}\right)\left(\frac{T}{20\,\text{mK}}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{1/2}\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right).$ (25)
+Thus, when $\exp\left(-(E_{f}-E_{c})^{2}/\Delta E^{2}\right)<10^{-1}$, the dominant contribution to the total heat capacity is from the semiconductor, i.e., $C_{s}+C_{e}\approx C_{s}$.
 However, note the condition in Eq.(14), which requires $\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)A\exp\left(-(E_{f}-E_{c})^{2}/\Delta E^{2}\right)$ not to be too small.
@@
 Therefore, it is likely that choosing the Fermi energy $E_{f}$ such that
-$10^{-3}\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)^{-1}A^{-1}<\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right)<10^{-2}$ (26)
+$10^{-3}\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)^{-1}A^{-1}<\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right)<10^{-1}$ (26)
 which ensures both that $C_{s}>C_{e}$ and that the microwave is effectively absorbed.
@@
 Now, when the sample absorbs the amplified microwave radiation, the rate of temperature increase is given by
-$\dot{T}=\frac{P_{ra}}{C_{s}(T)}\simeq\frac{7.2\,\text{mK}}{\text{s}}\,g_{\gamma}^{2}\left(\frac{20\,\text{mK}}{T}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}/\text{cm}^{3}}\right)$ (27)
-where we take the thickness $d=0.1\,\mu\text{m}$ and temperature $T=20\,\text{mK}$, with an external magnetic field $B_{t}=1.5\times 10^{5}\,\text{Gauss}$, applied parallel to the sample layers.
-( Even when smaller magnetic field $B_{t}=6\times 10^{4}$Gauss, $\dot{T}\simeq 0.032$mK/ms with lower temperature T=10T=10mK.
+$\dot{T}=\frac{P_{ra}}{C_{s}(T)}\simeq\frac{0.72\,\text{mK}}{\text{s}}\,g_{\gamma}^{2}\left(\frac{20\,\text{mK}}{T}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}/\text{cm}^{3}}\right)$ (27)
+where we take the thickness $d=1\,\mu\text{m}$ and temperature $T=20\,\text{mK}$, with an external magnetic field $B_{t}=1.5\times 10^{5}\,\text{Gauss}$, applied parallel to the sample layers.
+( Even when smaller magnetic field $B_{t}=10^{5}$Gauss, $\dot{T}\simeq 0.32$mK/s with lower temperature T=20T=20mK.
 )
@@
 where $\tau\equiv C_{s}/G$ is the time constant such that without microwave power ($P_{ra}=0$), the excess temperature decays as $\Delta T\propto\exp(-t/\tau)$ as $t\to\infty$.
-Thus, the final temperature increase due to axion-induced microwave absorption is approximately
-$\Delta T(t=\infty)\simeq$ $0.72$ $\,\text{mK}\left(\frac{\tau}{100\,\text{ms}}\right)\,g_{\gamma}^{2}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)$ (30)
-$\times$ $\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$
+[…]
+Thus, the temperature rise we observe with the observation time $t_{ob}(<\tau)$ is approximately
+$\Delta T(t=\infty)\simeq$ $0.72$ $\,\text{mK}\left(\frac{t_{ob}}{1\,\text{s}}\right)\,g_{\gamma}^{2}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)$ (30)
+$\times$ $\left(\frac{1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$
 The formula is applicable only for QCD axion model in which the coupling constant $g_{a\gamma\gamma}$ satisfies $g_{a\gamma\gamma}\simeq 39g_{\gamma}(10^{-14}\text{GeV}^{-1})(m_{a}/10^{-5}\text{eV})$.
 Using model independent coupling constant $g_{a\gamma\gamma}$, we represent the temperature increase in the following,
-$\Delta T(t=\infty)\simeq$ $5$ $\text{mK}\left(\frac{\tau}{100\,\text{ms}}\right)\,\Big(\frac{g_{a\gamma\gamma}}{10^{-14}\text{GeV}^{-1}}\Big)^{2}\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)$ (31)
-$\times$ $\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$
-We show in Fig.(3) the excluded region of $g_{a\gamma\gamma}$ and $m_{a}$ indicated by dashed line when we do not observe $\Delta T$ equal to or larger than $0.5\text{mK}(\tau/10\text{ms})$ in the case of time constant $\tau=10$ms.
-If the time constant $\tau$ is less than 1010ms, for instance, 11ms, the excluded region is above the same dashed line indicated with $\Delta T=0.05\text{mK}(\tau/1\text{ms})$ as the line in the figure when we do not observe $\Delta T\geq 0.05$mK.
+$\Delta T(t=\infty)\simeq$ $5$ $\text{mK}\left(\frac{t_{ob}}{1\,\text{s}}\right)\,\Big(\frac{g_{a\gamma\gamma}}{10^{-14}\text{GeV}^{-1}}\Big)^{2}\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)$ (31)
+$\times$ $\left(\frac{1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$
+We show in Fig.(3) the excluded region of $g_{a\gamma\gamma}$ and $m_{a}$ indicated by dashed line when we do not observe $\Delta T$ equal to or larger than $50\text{mK}(t_{ob}/1\text{s})$ with observation time $t_{ob}=1$s.
 In the figure, we have roughly depicted the excluded region by CAPPCAPP1 ; CAPP2 ; CAPP3 .
 We also show the lines associated with the models of QCD axion.
-Figure 3: When we do not observe $\Delta T$ larger than $0.5\text{mK}(\tau/10\text{ms})$ with time constant $\tau=10$ms, the value of $g_{a\gamma\gamma}$ above the dashed line denoted with $0.5\text{mK}(\tau/10\text{ms})$ is excluded.
+Figure 3: When we do not observe $\Delta T$ larger than $5\text{mK}(t_{ob}/1\text{s})$ with observation time $t_{ob}=1$s, the value of $g_{a\gamma\gamma}$ above the dashed line denoted with $0.5\text{mK}(\tau/10\text{ms})$ is excluded.
 We roughly show excluded region obtained by CAPP.
-Therefore, we find that a large time constant $\tau>10\,\text{ms}$ is preferable to observe a measurable temperature increase $\Delta T$.
-To achieve such a large time constant, it is needed to cover the semiconductor sample with a transparent, thin sheet that has both lower thermal conductivity and lower heat capacity than those of the sample semiconductor.
-Additionally, minimizing the heat capacity $C_{s}$ by using a very thin sample (e.g., $d\leq 0.1\,\mu\text{m}$) is advantageous.
-Importantly, the result $\Delta T$ does not depend on the surface area $S$ of the sample since both $P_{ra}\propto S$ and $C_{s}\propto S$; $\Delta T\propto P_{ra}/C_{s}$.
+Therefore, we find that a large time constant $\tau>1\,\text{s}$ is preferable to observe a measurable temperature increase $\Delta T$.
+To achieve such a large time constant, it is needed to minimize thermal conductance $G$ using possibly small copper lead wire and pedestal.
+In order to measure the temperature rise of the order of 11mK, we may adopt standard ruthenium oxide resistance thermometer, which is used for the experiment of quantum Hall effect even at low temperature $\sim 20$mK.
 To identify the resonance condition, we finely tune the cavity length $l$.
@@
 While the measurement time per step may vary depending on the detection method, this approach provides a highly efficient means of searching for axions.
-Our idea for the detection of dark matter axion is that we measure the increase of the temperature caused by the absorption of much weak microwaves generated by the axion.
-So it is favorable to use materials with much small heat capacity.
+To detect axion, we measure the increase of the temperature of material caused by the absorption of amplified but still much weak microwaves generated by the axion.
+So it is favorable to use material with much small heat capacity and small thermal conductance G of the material connecting to heat bath.
 In the paper we have proposed two dimensional electrons forming quantum Hall states as such a material.
@@
 […]
-Although the method for measuring the temperature rise is beyond the scope of this work, the predicted increase offers a direct and sensitive means for detecting the axion and determining its mass.
+Although the detailed method how we measure the temperature rise is beyond the scope of this work, the predicted increase offers a direct and sensitive means for detecting the axion and determining its mass.
 The author expresses thanks to A. Sawada for useful comments.